# Flexible Process Variant Binding in Information Systems with Software Product Line Engineering

### v1

@@ 5.1 Properties
In line with Requirements R1-R6 (cf. Section 4.1), the proof of concept needs to provide the following properties. In addition, the development phases of a PAIS Product Line shall be tool-supported similar to SPL Engineering, which results in P7.
@@ 1. P1:
    PAIS Products are derivable from a PAIS Product Line. All derivable PAIS Products include the mandatory activity implementations, but may differ in the selection of alternative activity implementations. The proof of concept has to show how PAIS Products can be derived from a PAIS Product Line from a technical perspective, i.e. what variability mechanism is used.
@@ 2. P2:
    Multiple implementation of an activity might be selected and executed. A suitable mechanism needs to be chosen that allows to execute multiple implementations for one activity.
@@ 3. P3:
    A variability mechanisms needs to be used to compose the process data structure during PAIS Product derivation to meet the organisations’ individuals needs.
@@ 5.2 Data Structure Composition
In our previous work, the build tool Apache Maven was used to compose the artefacts during PAIS Product derivation. The data structure was the same among all PAIS Products and the activity implementations were designed as independent logical units and were compiled JARs (Java applications) that were conditionally included when the implementation was selected based on Apache Maven profiles. That is, Apache Maven was used to compose fully compiled artefacts. Note that this worked, as these logical units were self-contained and could be put together like building blocks. In contrast, when the process data structure needs to be customizable (cf. Example 4.4), there are changes on the class level, i.e. before its compilation the process data structure needs to be customised. Therefore, instead of composing JARs, Java classes need to be composed. To compose Java classes, both FeatureHouse [4] and the AHEAD tools [7] may be used. As FeatureHouse does not require any additional key words like refines, FeatureHouse is used to compose the Java classes of the process data structure in the proof of concept.
@@ 5.3 Multiple Implementations of an Activity
Specific activities shall constitute a variation point, i.e. one or more implementations shall be selectable for these activities. An implementation may be a service task, a user task, or a call activity, which calls a subprocess. During PAIS Product derivation, the core process model needs to be composed with the implementations. In this context, the implementations may be considered as refinements of the base process model. The language Xak [3] can be used to compose the process models. Note that the latter are modelled in terms of BPMN 2.0, which has an XML representation.
However, when using Xak, the refinements (i.e. the implementations) are not schema-compliant, i.e. the files in which an activity is refined (which are process models as well) cannot be edited in or displayed by BPMN 2.0 modelling tools. As one advantage, BPMN 2.0 has both an XML and a graphical representation of the process model. The graphical representation serves as a means of communicating process models between domain and IT experts. If the process models serving as implementations are not schema-compliant, the advantage of BPMN 2.0, i.e. the graphical representation, does not come into effect. Furthermore, when selecting lots of implementations these implementations need to be included in the core process model and the elements need to be rearranged in order to make room for the additional elements of the implementations. Although works such as [46] provide means for automatically layouting process models, with increasing number of included implementations the process model gets unclear and confusing. Therefore, Xak is not used to compose the base process model and its implementations. Instead, multi-instance call activities are used, which call the implementations as subprocesses. The implementation (i.e. automated task or user task) are embedded in a process model, which is then called by the variation point activity. Figure 8 shows a simple process with a variation point activity which corresponds to a multi-instance call activity that calls two implementations. The implementations, in turn, are simple processes consisting of an activity, which is either an automated task or a user task. As the variation point is a multi-instance activity it may invoke both implementations in parallel allowing for the execution of multiple implementations at the same time.
@@
Figure 9 shows the process model of Example 4.1 for the proof of concept. Activities Check application and Notify craftsperson are variation points, for which one or more implementations may be selected. Therefore these activities are multi-instance call activities, i.e. they call the corresponding implementations as subprocesses. A detailed description on how the implementations are bound and called by the core process model is provided in Section 5.4.
@@ 5.4 Static and Dynamic Feature Binding
The plugin mechanism allowing for static and dynamic feature binding in PAIS Product Lines is illustrated in Figure 10. During compilation, implementation process models may be included or excluded (static binding, cf. Figure 10 (1)). Each implementation process model has a unique identifier (ID), which is used to unambiguously identify the former. When an implementation process model is included during compilation, it is registered as a plugin using its ID with the corresponding variation point activity (cf. Figure 10 (2)). The excluded implementations cannot be registered as plugins and are not available at runtime. Implementation Automatic2 in Figure 10 is not included during compilation and consequently not registered as plugin. At start time, developers and, at runtime, users may deselect one or more implementations that have been registered for a variation point activity by setting parameters. Figure 10 (3) shows a user form that can be used to start the process by selecting/excluding specific plugins (i.e. implementations). Before a variation point activity calls the registered implementation process models it is checked whether one of the latter is excluded via a parameter (cf. Figure 10 (4)). Only implementation process models are invoked that are not excluded via parameters (dynamic binding).
@@
From a technical point of view, for static binding, FeatureHouse is used for composing the activity implementations (like for process data structure composition). The core process is bundled with the implementation process models and the implementing logic, namely HTML code for user tasks and Java code for automated tasks. After bundling the files in one folder, the PAIS Product can be compiled. At both start time and runtime, parameters may prevent certain implementations from being selected and executed. In conditional statements in software programs, frequently parameters are used to decide whether or not a specific feature is executed based on the value of the parameter. By providing values, parameters can be exploited as variability mechanism for feature binding at runtime. Note that the use of parameters as variability mechanisms are well known in literature [2][26].
Examples 10 and 5.4 are used to demonstrate how static and dynamic binding of activity implementations works by registering them as plugins. {customExample} (Feature Binding Time in PAIS Product Lines): Activity Notify craftsperson in Figure 9 of business process special parking permit for craftsperson is used to notify the applicant when his/her application was rejected. This activity is a variation point for which multiple implementations are available. The applicant may be notified via sms, e-mail, or manually by a clerk. At compile time and start time, implementations may be selected by a developer. At runtime, the applicant can decide how he/she wants to be notified from the available notification means that have been included by a developer.
@@ 12 Map<String, String> pluginLabelsById = new HashMap<String, String>();
23 Map<String, String> pluginProcessIdsByPluginId = new HashMap<String, String>();
@@
In summary, for every activity that constitutes a variation point there is a plugin mechanism. Every implementation needs to implement the interface of the activity it wants to register with. At compile time (i.e. static binding) activity implementations may be included or excluded using FeatureHouse. The activity implementations consists of the implementation process model and the associated code base (Java code for automated tasks and html code for user tasks). At start time, command line parameters, and, at runtime, user input may be used to deselect some activity implementations (dynamic binding).
@@ 5.5 Tool Support
To support the phases of SPL Engineering with a tool, FeatureIDE is used in the proof of concept. A feature model (cf. Figure 14) visualises the variability of the PAIS Product Line during domain analysis. Note that the feature model in Figure 14 is structured from left to right such that it fits on the page. FeatureIDE captures the constraints in text form below the feature model as opposed to arrows used in other approaches.
@@
During requirements analysis, FeatureIDE supports the developer by providing a configuration editor (cf. Figure 15). In this configuration editor the developer can select the desired activity implementations that shall be included in the PAIS Product. For implementing the variability of the PAIS Product Line, FeatureHouse was chosen, which is supported by FeatureIDE. When deriving a PAIS Product, FeatureIDE uses the the selection from the configuration editor as input (Figure 15) to compose the variable activity implementations.
@@ 6 Discussion
The presented approach has shown how to derive products from PAIS Product Lines with varying process data structures, which use multiple implementations for one activity, and that can bind features at compile time, start time, and runtime. In the following, the approach is assessed in respect to whether it meets the requirements set out in Section 4.1. Furthermore, the proof of concept is tested to see whether it provides the listed properties. Finally, limitations are disclosed.
@@ 6.1 Requirements
A PAIS Product Line corresponds to a set of common core artefacts from which PAIS Products may be derived. Hence, Requirement R1 is satified. Multiple implementations can be selected for one activity. When the activity may write to the process data structure, aggregation code needs to be included in order to prevent the implementations from overwriting each others data. Consequently, Requirement R2 is met. Activity implementations may require adaptions to the process data structure. When selecting a specific activity implementation the required part of the data structure needs to be selected as well. Therefore, Requirement R3 is satisfied. Requirement R4 is met as a process feature model represents the variability of a PAIS Product Line with the corresponding constraints. Unintended feature interaction can occur only when the activity implementations write to the process data structure. As aggregation code prevents implementations from overwriting each others data Requirement R5 is satisfied. At runtime, activity implementations are selected per process instance. Once they are selected, they cannot be unselected anymore. Furthermore, activity implementations can be run independently and have no technical dependency among each other. Consequently, activity implementations do not share system processes or data that might be corrupt. In addition, data does not get lost due to terminated activity implementations. Hence, Requirement R6 is satisfied as well.
@@ 6.2 Properties
The PAIS Product Line of the proof of concept uses FeatureHouse to include optional activity implementations. When including an activity implementation it registers as plugin with the corresponding activity. Consequently, the proof of concept provides Properties P1 and P2. FeatureHouse is used to compose the process data structure as well. Hence, Property P3 is provided. As a process feature model outlines the variability of the PAIS Product Line Property, P4 is provided. Aggregation code is implemented as a Camunda variable mapper class, which is conditionally included using FeatureHouse. Therefore, Property P5 is provided. Static and dynamic feature binding is supported by using a mix of FeatureHouse and parameters. Thus, Property P6 is provided. Since FeatureIDE is used as tool support in the proof of concept Property P7 is provided.

### v2

@@ 5.1 Properties
In line with Requirements R1-R6 (cf. Section 4.1), the proof-of-concept needs to provide the following properties. In addition, the development phases of a PAIS Product Line shall be tool-supported similar to SPL Engineering, which results in P7.
@@ 1. P1:
    PAIS Products are derivable from a PAIS Product Line. All derivable PAIS Products include the mandatory activity implementations, but may differ in the selection of alternative activity implementations. The proof-of-concept has to show how PAIS Products can be derived from a PAIS Product Line from a technical perspective, i.e. what variability mechanism is used.
@@ 2. P2:
    Multiple implementations of an activity might be selected and executed. A suitable mechanism needs to be chosen that allows to execute multiple implementations for one activity.
@@ 3. P3:
    A variability mechanism needs to be used to compose the process data structure during PAIS Product derivation to meet the organisations’ individuals needs.
@@ 5.2 Data Structure Composition
In our previous work, the build tool Apache Maven was used to compose the artefacts during PAIS Product derivation. The data structure was the same among all PAIS Products and the activity implementations were designed as independent logical units and were compiled JARs (Java applications) that were conditionally included when the implementation was selected based on Apache Maven profiles. That is, Apache Maven was used to compose fully compiled artefacts. Note that this worked, as these logical units were self-contained and could be put together like building blocks. In contrast, when the process data structure needs to be customizable (cf. Example 4.4), there are changes on the class level, i.e. before its compilation the process data structure needs to be customised. Therefore, instead of composing JARs, Java classes need to be composed. To compose Java classes, both FeatureHouse (Apel et al., 2009) and the AHEAD tools (Batory et al., 2004) may be used. As FeatureHouse does not require any additional key words like refines, FeatureHouse is used to compose the Java classes of the process data structure in the proof-of-concept.
@@ 5.3 Multiple Implementations of an Activity
Specific activities shall constitute a variation point, i.e. one or more implementations shall be selectable for these activities. An implementation may be a service task, a user task, or a call activity, which calls a subprocess. During PAIS Product derivation, the core process model needs to be composed with the implementations. In this context, the implementations may be considered as refinements of the base process model. The language Xak (Anfurrutia et al., 2007) can be used to compose the process models. Note that the latter are modelled in terms of BPMN 2.0, which has an XML representation.
However, when using Xak, the refinements (i.e. the implementations) are not schema-compliant, i.e. the files in which an activity is refined (which are process models as well) cannot be edited in or displayed by BPMN 2.0 modelling tools. As one advantage, BPMN 2.0 has both an XML and a graphical representation of the process model. The graphical representation serves as a means of communicating process models between domain and IT experts. If the process models serving as implementations are not schema-compliant, the advantage of BPMN 2.0, i.e. the graphical representation, does not come into effect. Furthermore, when selecting lots of implementations these implementations need to be included in the core process model and the elements need to be rearranged in order to make room for the additional elements of the implementations. Although works such as Scholz and Lübke (2019) provide means for automatically layouting process models, with increasing number of included implementations the process model gets unclear and confusing. Therefore, Xak is not used to compose the base process model and its implementations. Instead, multi-instance call activities are used, which call the implementations as subprocesses. The implementations (i.e. automated task or user task) are embedded in a process model, which is then called by the variation point activity. Figure 8 shows a simple process with a variation point activity which corresponds to a multi-instance call activity that calls two implementations (Note that the two dashed arrows are not part of BPMN but rather illustrate how a call activity calls subprocesses). The implementations, in turn, are simple processes consisting of an activity, which is either an automated task or a user task. As the variation point is a multi-instance activity it may invoke both implementations in parallel allowing for the execution of multiple implementations at the same time.
@@
Figure 9 shows the process model of Example 4.1 for the proof-of-concept. Activities Check application and Notify craftsperson are variation points, for which one or more implementations may be selected. Therefore, these activities are multi-instance call activities, i.e. they call the corresponding implementations as subprocesses. A detailed description on how the implementations are bound and called by the core process model is provided in Section 5.4. Note that activity Apply for parking permit is removed in this process model. Due to technical characteristics of Camunda Platform, the application form to apply for the parking permit is used to start the process and, hence, is not contained in the process model.
@@ 5.4 Static and Dynamic Feature Binding
The plugin mechanism allowing for static and dynamic feature binding in PAIS Product Lines is illustrated in Figure 10. Figure 10 contains a process model with a variation point as well as multiple implementation process models that can be invoked by this variation point. The processes are modelled using BPMN whereas solid boxes show in which components the process models are organised. The dashed box and arrows describe the steps to achieve variability from compilation to runtime (Note that the boxes and arrows are not part of the BPMN specification). During compilation, implementation process models may be included or excluded (static binding, cf. Figure 10 (1)). Each implementation process model has a unique identifier (ID), which is used to unambiguously identify the former. When an implementation process model is included during compilation, it is registered as a plugin using its ID with the corresponding variation point activity (cf. Figure 10 (2)). The excluded implementations cannot be registered as plugins and are not available at runtime. Implementation Automatic2 in Figure 10 is not included during compilation and consequently not registered as plugin. At start time, developers and, at runtime, users may deselect one or more implementations that have been registered for a variation point activity by setting parameters. Figure 10 (3) shows a user form that can be used to start the process by selecting/excluding specific plugins (i.e. implementations). Before a variation point activity calls the registered implementation process models it is checked whether one of the latter is excluded via a parameter (cf. Figure 10 (4)). Only implementation process models are invoked that are not excluded via parameters (dynamic binding).
@@
From a technical point of view, for static binding, FeatureHouse is used for composing the activity implementations (like for process data structure composition). The core process is bundled with the implementation process models and the implementing logic, namely HTML code for user tasks and Java code for automated tasks as well as configuration files (i.e. application.properties files for Spring Boot). As described in Section 2.4, FeatureHouse represents artefacts including their hierarchical structure as a tree and merges them. That is, each feature corresponds to a folder containing the software artefacts within their hierarchical structure, i.e. the Java files in packages but also the process models and configuration files in their respective folders. During composition, FeatureHouse merges the folders of the selected features taking into account the hierarchies within the folders (i.e. packages and folders). In the proof-of-concept, each feature folder consists of the packages for the Java classes, a folder for the implementation process models, and when necessary a folder for the application.properties (i.e. the files in Spring Boot containing configuration, e.g. URLs to web services, credentials). While Java classes are composed on the class level, i.e. the fields within a Java class split up across multiple feature folders are merged, other artefacts like process models and application.properties files are merge by copying the artefacts into the same folder. By default, Camunda Platform deploys all BPMN process files in the classpath resources folder. To apply the configurations, we have configured Spring Boot in that all application.properties whose path in the classpath corresponds to a defined pattern are loaded. After composition, the various artefacts (i.e. Java classes, process models, and application.properties) are arranged in the Maven standard structure so that the PAIS Product can be compiled natively with Maven. At both start time and runtime, parameters may prevent certain implementations from being selected and executed. In conditional statements in software programs, frequently parameters are used to decide whether or not a specific feature is executed based on the value of the parameter. By providing values, parameters can be exploited as variability mechanism for feature binding at runtime. Note that the use of parameters as variability mechanisms is well known in literature (Anastasopoules and Gacek, 2001; Jacobson et al., 1997).
Examples 10 and 5.4 are used to demonstrate how static and dynamic binding of activity implementations works by registering them as plugins. {customExample} (Feature Binding Time in PAIS Product Lines): Activity Notify craftsperson in Figure 9 of business process special parking permit for craftsperson is used to notify the applicant when his/her application was rejected. This activity is a variation point for which multiple implementations are available. The applicant may be notified via SMS, e-mail, or manually by a clerk. At compile time and start time, implementations may be selected by a developer. At runtime, the applicants can decide how they want to be notified from the available notification means that have been included by a developer.
@@ 12 Map<String, String> pluginLabelsById = new HashMap<String, String>();
23 Map<String, String> pluginProcessIdsByPluginId = new HashMap<>();
@@
From a technical perspective, the variable mapper (i.e. aggregation code) accepts the data fields justified from each subprocesses and writes them to a list in the core process model. Then, the variable mapper continuously iterates over the list and checks whether all values are true and therefore the application as a whole is considered justified. Consequently, there are no race conditions as the order in which the subprocess complete does not matter. In any case, the results of the subprocesses are added to a list and subsequently evaluated. Subprocesses do not overwrite each others results.
Besides race conditions, during feature binding, there might occur other technical problems. 1) Data integrity might be impacted when executing multiple activity implementations and consequently writing process data concurrently (data corruption). Furthermore, 2) activity implementations might be cancelled or unbound due to failures during execution (e.g. an activity implementation tries to call a web service that is not available). These are typical problems not only when digitising business processes variants but also when digitising regular business processes for which most WfMSs provide mechanisms. As mentioned before, Camunda Platform was not customised and call activities and subprocesses are used for feature binding at runtime. Consequently, we are able to solely rely on Camunda Platform’s error handling and resilience features. In the following, it is shown how we use Camunda Platform to solve concurrency problems and errors during the execution of activity implementation.
1) In general, if multiple threads write concurrently data to a process, e.g. two activities complete at the same time, Camunda Platform throws an OptimisticLockingExceptions, rolls back one of the transactions and retries [^9] https://docs.camunda.org/manual/7.22/user-guide/process-engine/transactions-in-processes/#handling-optimistic-locking-exceptions. That is, one thread writes the data whereas the other retries writing the data a moment later. When configured accordingly and applied to our approach, the activity implementations realised as subprocesses invoked by a call activity are each executed in an own thread, which results in concurrency. Consequently, Camunda Platform manages the concurrent access to process data and hence data cannot be corrupted.
2) In general, if the execution of an activity fails regardless of the underlying error (e.g. a web service is not available), Camunda Platform uses a retry mechanism in order to continuously try executing the activity. The repetition intervals and the maximum number of repetitions can be configured in Camunda Platform. After the maximum number of repetitions has been reached, Camunda Platform raises an incident in the management interface, which is called Cockpit. Then, an administrator can decide how to proceed. In most cases, when the error occurred due to a non-available web service, the execution can just be resumed. However, the administrator has other options as well, e.g. change process data of the running process instance. Note that during a rollback, data gathered since the last transaction boundary is lost. Transaction boundaries and hence how far back to roll can explicitly be specified in business processes executed in Camunda Platform [^10] https://docs.camunda.org/manual/7.22/user-guide/process-engine/transactions-in-processes/#transaction-boundaries. However, a failure during the execution of one subprocess of a multi-instance call activity does not impact the execution of others. This retry mechanism applies to our usage of multi-instance call activity for feature binding as well [^11] https://docs.camunda.org/manual/7.22/user-guide/process-engine/the-job-executor/#use-a-custom-job-retry-configuration-for-multi-instance-activities. Consequently, unless an administrator intentionally cancels a subprocess, features/activity implementations are not unbound during execution.
@@
In summary, for every activity that constitutes a variation point there is a plugin mechanism. Every implementation needs to implement the interface of the activity it wants to register with. At compile time (i.e. static binding) activity implementations may be included or excluded using FeatureHouse. The activity implementations consist of the implementation process model and the associated code base (Java code for automated tasks and html code for user tasks). At start time, command line parameters, and, at runtime, user input may be used to deselect some activity implementations (dynamic binding).
@@ 5.5 Tool Support
To support the phases of SPL Engineering with a tool, FeatureIDE is used in the proof-of-concept. A feature model (cf. Figure 14) visualises the variability of the PAIS Product Line during domain analysis. Note that the feature model in Figure 14 is structured from left to right such that it fits on the page. FeatureIDE captures the constraints in text form below the feature model as opposed to arrows used in other approaches.
@@
During requirements analysis, FeatureIDE supports the developer by providing a configuration editor (cf. Figure 15). In this configuration editor the developer can select the desired activity implementations that shall be included in the PAIS Product. For implementing the variability of the PAIS Product Line, FeatureHouse was chosen, which is supported by FeatureIDE. When deriving a PAIS Product, FeatureIDE uses the selection from the configuration editor as input (Figure 15) to compose the variable activity implementations.
@@ 6 Discussion
The presented approach has shown how to derive products from PAIS Product Lines with varying process data structures, which use multiple implementations for one activity, and that can bind features at compile time, start time, and runtime. In the following, the approach is assessed in respect to whether it meets the requirements set out in Section 4.1. Furthermore, the proof-of-concept is tested to see whether it provides the listed properties. Finally, limitations are disclosed.
@@ 6.1 Requirements
A PAIS Product Line corresponds to a set of common core artefacts from which PAIS Products may be derived. Hence, Requirement R1 is satisfied. Multiple implementations can be selected for one activity. When the activity may write to the process data structure, aggregation code needs to be included in order to prevent the implementations from overwriting each others’ data. Consequently, Requirement R2 is met. Activity implementations may require adaptions to the process data structure. When selecting a specific activity implementation, the required part of the data structure needs to be selected as well. Therefore, Requirement R3 is satisfied. Requirement R4 is met as a process feature model represents the variability of a PAIS Product Line with the corresponding constraints. Unintended feature interaction can occur only when the activity implementations write to the process data structure. As aggregation code prevents implementations from overwriting each others’ data Requirement R5 is satisfied. At start time, activity implementations are selected per derived PAIS, which are not unbound while the PAIS is running. Consequently, data cannot be lost. At runtime, activity implementations are selected per process instance. Once they are selected, they cannot be unselected anymore. Furthermore, activity implementations can be run independently and have no technical dependency among each other. Consequently, activity implementations do not share system processes or data that might be corrupt. In addition, data does not get lost due to terminated activity implementations. Hence, Requirement R6 is satisfied as well.
@@ 6.2 Properties
The PAIS Product Line of the proof-of-concept uses FeatureHouse to include optional activity implementations. When including an activity implementation, it registers as plugin with the corresponding activity. Consequently, the proof-of-concept provides Properties P1 and P2. FeatureHouse is used to compose the process data structure as well. Hence, Property P3 is provided. As a process feature model outlines the variability of the PAIS Product Line Property, P4 is provided. Aggregation code is implemented as a Camunda variable mapper class, which is conditionally included using FeatureHouse. Therefore, Property P5 is provided. Static and dynamic feature binding is supported by using a mix of FeatureHouse and parameters. Thus, Property P6 is provided. Since FeatureIDE is used as tool support in the proof-of-concept Property P7 is provided.
6.3 Scalability
While the proof-of-concept is based on real-world requirements, we acknowledge the fact that there are business processes in German municipalities and industry with more activities and more implementation options. Therefore, the scalability of the technical setup of the proof-of-concept is discussed in the following.
The phases of domain analysis, domain implementation, requirements analysis, software generation (i.e. phases of SPL Engineering), and runtime (i.e. execution of a derived PAIS) need to be distinguished. For the proof-of-concept, the modelling tool and the configuration editor of FeatureIDE were used for domain analysis and requirements analysis, respectively. FeatureHouse, which is integrated into FeatureIDE, was used for domain implementation and software generation. Camunda Platform is embedded as Java library in the proof-of-concept and executes the process models at runtime. Consequently, in the following, it is shown how well FeatureIDE, FeatureHouse, and Camunda Platform scale.
FeatureIDE is used in industry, which is the result of a survey aimed at gaining insights about feature modelling and feature modelling tools, targeting exclusively practitioners with industrial experience (Berger et al., 2013). In (Lettner et al., 2015), FeatureIDE was extended to be able to create hierarchical feature collections represented by feature models and define dependencies among features of the different collections. Then, FeatureIDE was used to model two systems, which both have around 400 features of which some are mandatory, optional, and alternative. Similarly, Linsbauer et al. (2022) extended FeatureIDE to allow for partial configuration enabling splitting the product configuration among different groups of stakeholders so that the product is gradually configured. This extension is now part of the official distribution of FeatureIDE, which we used for our proof-of-concept. The authors evaluated their approach with five case studies of which one uses FeatureHouse as variability mechanism (i.e. the variability mechanism we use in our proof-of-concept) and contains 119 features.
Apel et al. (2013) have conducted 50 case studies in which they applied FeatureHouse to compose software artefacts to software systems. The case studies include software systems written in various languages (e.g. Java) consisting of up to 99 features and approximately 64 thousand lines of code. The authors conclude that FeatureHouse scales well with the number of features and lines of code.
After the artefacts have been composed using FeatureIDE and FeatureHouse, the proof-of-concept is natively compiled as Java application and Camunda Platform is used as an embedded Java library to execute the process models. Camunda Platform has been benchmarked among other WfMSs (Skouradaki et al., 2016). The results show that the performance of Camunda Platform depends on the used workflow pattern. However, generally speaking, Camunda Platform is able to execute in average more than thousand process instances per second when using a mix of workflow patterns. The authors conclude that Camunda Platform performed well in comparison to the other investigated WfMSs. Based on our own experience over the last six years working with Camunda Platform in industry and public sector, we can confirm that Camunda Platform is able to execute several tens of thousand processes concurrently.
In summary, FeatureIDE and FeatureHouse have been investigated in numerous case studies, in which they scaled well separately as well as in combination. After the PAIS Product is derived (i.e. the artefacts are composed), the proof-of-concept is natively compiled as plain Java application, which includes Camunda Platform as WfMS. Camunda Platform has proven suitable to execute process models in scaled settings. As FeatureIDE and FeatureHouse are only used to compose the artefacts for compilation, they cannot influence the runtime. Consequently, interferences between FeatureIDE, FeatureHouse on the one hand, and Camunda Platform on the other are not to be expected. In conclusion, we are confident that the technical setup of our proof-of-concept as a whole scales well for business processes with lots of activities and various optional and alternative activity implementations.